# HateSieve: A Contrastive Learning Framework for Detecting and Segmenting Hateful Content in Multimodal Memes

### v1

@@ 4.1 Setup
Our dataset generation for the triplet model leverages the HatefulMemes Kiela et al. (2020) and MMHS150k Gomez et al. (2020) datasets. We evaluated our framework by validating the classification and segmentation effects using the HatefulMemes test-unseen dataset. This dataset, provided primarily by Facebook Research, comprises 10,000 annotated meme images. Each meme integrates text and an image to challenge the model’s proficiency in detecting hate speech within memes. The MMHS150k dataset contains 150,000 tweets, each featuring an image and text pair, compiled using 51 terms from Hatebase. These tweets were collected between September 2018 and February 2019, specifically to analyze hate speech expressions on social media. Our training dataset for contrastive learning consists of 8,500 entries from the HatefulMemes training dataset and 33,844 hateful data samples from MMHS150k by our contrastive meme generator.
@@
In the classification evaluation for our HateSieve, our baseline models that we compare to are as follows:
  • LMMs: The set includes GPT-4V Achiam et al. (2023), CogVLM Wang et al. (2023), LLava-1.5 Liu et al. (2023), InstructBLIP Dai et al. (2024), MiniGPT-4 Zhu et al. (2023), Qwen-VL Bai et al. (2023), OpenFlamingo Awadalla et al. (2023), MMGPT Gong et al. (2023), and MinGPT-v2 Chen et al. (2023) for zero-shot inference, along with LLava-1.5, InstructBLIP, and BLIP2 employing Supervised Fine-Tuning with QLora Dettmers et al. (2024).
  • CLIP-based method: It encompasses the original CLIP model along with its variants, HateCLIPer and MOMENTA.
We provide detailed descriptions of the methods we implemented for these experiments in Appendix A.
Evaluating the segmentation capabilities of HateSieve is challenging due to the absence of pixel-level and token-level annotations. Consequently, we sampled 100 memes from the HatefulMemes dataset and manually compared the segmentation performance of our HateSieve model against LMMs equipped with vision prompt generators such as InstructBLIP. We involved three human annotators in the process, and decisions were made based on a majority vote.
Implementation Details
Using the Contrastive Meme Generator, we produced a total of 42,344 triplet pairs. During the pre-training and fine-tuning phases, we employed the clip-vit-base-patch32 as our backbone for the image-text encoder and froze all the CLIP parameters. Our newly introduced Image-Text Alignment module comprises 6 layers of self-attention blocks. Additionally, we incorporated a two-layer MLP as a decoder for classification fine-tuning.
4.2 Results on HatefulMeme Classification
Models Acc. F1 # t.para.
@@ 4.2 Results on HatefulMeme Classification
GPT-4V (-) 71.70 71.28 -
CogVLM (17B) 61.50 60.03 -
LLaVA-1.5 (13B) 65.20 61.40 -
InstructBLIP (13B) 58.25 57.42 -
MiniGPT-4 (13B) 58.20 39.98 -
Qwen-VL (10B) 64.00 56.42 -
OpenFlamingo (9B) 58.65 51.78 -
MMGPT (9B) 37.50 27.28 -
MiniGPT-v2 (7B) 57.35 57.27 -
BLIP2 (6.7B) 56.34 55.29 -
Supervised Fine-Tuning
InstructBLIP (13B) 63.55 59.34 65.72M
LLaVA-1.5 (13B) 66.34 63.28 65.72M
BLIP2 (6.7B) 62.85 56.43 33.35M
@@
MOMENTA (434M) 73.34 70.02 7.73M
HateSieve-ours (155M) 73.45 71.64 3.61M
-w/o Constrastive Learning 69.38 65.62 3.61M
Table 1: Model Performance Comparison. The bold scores highlight the best performance, while the underline scores signify the second-best performance. ’Acc.’ and ’F1’ represent the classification accuracy and macro-F1 score, respectively. ’#t.param’ denotes the total number of trainable parameters within the model.
As shown in table 1, we assessed the efficacy of various LMMs and CLIP-based methods in zero-shot inference and SFT settings. Overall, except for GPT-4V, other open-source models demonstrated limited capabilities under zero-shot conditions, revealing that while current LMMs exhibit multimodal recognition and understanding abilities from extensive pre-training, they still fall short in effectively recognizing hateful memes that demand more complex semantic reasoning. In the supervised fine-tuning setting, the performance improvements of LMMs, even for those fine-tuned with QLora, were minimal. For a detailed visual analysis, please see section 4.3.
@@
4.3 Results on Hate Component Segmentation
Figure 7: Hate Component Segmentation Analysis: Comparing InstructBLIP, CLIP+ITA, and our HateSieve.
We compare InstructBLIP after SFT on the HatefulMeme dataset, CLIP+ITA which is directly fine-tuned with HatefulMeme, and our HateSieve model. For InstructBLIP, we prompt the model as follows: “Please examine the provided meme, which includes an [image] and accompanying [text]. Determine if the content can be considered hateful. If you conclude that the meme is hateful, identify and list the specific keywords or phrases in the text.” This allows us to identify the text tokens that InstructBLIP considers ambiguous. Regarding image segmentation, we adhere to Li et al. approach, mapping the query corresponding to the Q-Former in InstructBLIP with the image’s cross-attention map using bilinear interpolation.
InstructBLIP CLIP+ITA HateSieve
# of Valid. 21 13 57
Table 2: Manual evaluation of the segmentation on 100 random samples. ’# of Valid’ indicates the count of heatmaps confirmed as valid in the manual verification.
As illustrated in the figure 7, the attention heatmap generated by InstructBLIP converges to a very small area, resulting in ineffective subsequent segmentation. Although InstructBLIP classifies this meme as hateful, it indicates that the model does not fully utilize the image information in its analysis. In contrast, the attention map of CLIP+ITA, which is solely fine-tuned on HatefulMeme, is overly dispersed across the image and fails to concentrate on the relevant characters or scenes. On the other hand, our contrastively pre-trained HateSieve effectively focuses on specific characters in the image and correlates them with the keywords in the text. This demonstrates that our contrastive learning pre-training enables HateSieve to correctly link specific components of the image and of the text within the meme.
Due to the lack of precise image and text annotations, we randomly sampled 100 groups of memes from HatefulMeme and manually evaluated the accuracy of the images rendered by the attention heatmap and the corresponding text. As shown in Table 2, our HateSieve significantly outperforms InstructBLIP and CLIP+ITA in segmentation effectiveness, validating the positive correlation of contrastive learning with image-text associations.
@@ 5 Ablation Study
Figure 8: Impact of ITA Parameter Scale and Triplet Data Scale on Model Performance.
Figure 9: Impact of using different pre-training strategies.
@@
We investigated the influence of self-attention layers within the ITA on the classification efficacy of HateSieve. As depicted in Figure 8, there is a clear trend where the classification performance of HateSieve improves when increasing the number of layers. However, this improvement plateaus and subsequently declines once the layer count surpasses 6, as evidenced by a noticeable drop in the F1 score.
@@
We investigated the impact of the triplet dataset scale used during the pre-training stage with contrastive learning. As illustrated in Figure 8, we analyzed the effects on the classification performance of HateSieve when pre-trained with 0% (no pre-training), 25%, 50%, 75%, and 100% of the triplet dataset. The results clearly show that an increase in pre-training data enhances the classification capabilities of HateSieve. This underscores the critical role of our triplet dataset in the effectiveness of contrastive learning during pre-training.
@@
We also investigated the impact of different pre-training strategies on the classification performance of HateSieve. We compared the following approaches:
  • In-domain pre-training, which involves using the HatefulMeme training set to directly sample similar hateful version memes, rather than sampling $\{I^{-},T^{-}\}$ from the MMHS150k dataset during the triplet dataset generation process.
  • CLS (Classification-only Strategy): In this approach, we substituted the contrastive learning pre-training with direct classification. We utilized the same volume of training samples, comprising 8,500 entries from the HatefulMeme dataset and 33,844 sets of multimodal hate speech data from MMHS150k.
As shown in Figure 9, the model pre-trained in-domain outperformed the CLS pre-training, confirming the superior effectiveness of contrastive learning. However, it underperformed compared to other HateSieve resources, We speculate that this lower performance is due to the limited variety in the HatefulMeme dataset and unseen test data. Sole reliance on the in-domain dataset didn’t substantially improve generalization.
Meanwhile, although the CLS-only pre-trained model improved upon the baseline HateSieve, as shown in Table 1, the significant computational expense did not match the modest performance gains, making it inefficient.
6 Segmentation Results
@@ 7 Conclusion
We developed HateSieve, a framework for classifying and segmenting hateful memes. Our experiments show that contrastive learning on a custom triplet dataset improves classification and achieves effective segmentation.
@@
Our current work faces several limitations, including the broad semantic scope of our contrastive meme generator. We plan to refine this to enhance contrastive learning in future iterations. Additionally, achieving high accuracy for image segmentation in HateSieve is challenging. The attention maps are restricted to the image-patch level, and attempts to refine these to pixel-level detail via linear interpolation introduce biases and do not inherently optimize for segmentation accuracy. We aim to expand these capabilities and also to conduct further experiments on a more extensive array of multimodal content related to hate speech in future work.
@@ Appendix A Appendix
A.1 LMMs Hyperparameters
@@
When implementing supervised fine-tuning on LMMs, we consistently utilized the QLora framework by integrating a set of trainable parameters ($d$=64) into both the query and key components of the Q-Former. This adjustment was applied to joint LLM architectures such as OPT-6.7b for BLIP2 and Vicuna-7b for InstructBLIP, while the original parameters were kept frozen. We established a constant dropout rate of 0.05 and set the hyperparameter $\alpha$ at 256. The fine-tuning was conducted with a learning rate of $55e-$ 5 and a batch size of 8.

### v2

@@ 4.1 Setup
To generate our triplet dataset, we utilized the HatefulMemes Kiela et al. (2020) and MMHS150k Gomez et al. (2020) datasets. For contrastive learning training, we incorporated 8,500 entries from the HatefulMemes training set and 33,844 hateful memes sampled from MMHS150k using our contrastive meme generator. For classification fine-tuning, we trained and evaluated our framework’s performance on the HatefulMemes test-unseen category, as well as on the Harm-C and Harm-P datasets Pramanick et al. (2021), employing a binary classification setting. Additionally, we assessed the effectiveness of our segmentation approach on the HatefulMemes dataset. Details of the dataset are in Appendix A.2.
@@
We compare our HateSieve framework against the following baseline models for classification task:
  • LMMs: We evaluate GPT-4V Achiam et al. (2023), CogVLM Wang et al. (2023), LLaVA-1.5 Liu et al. (2023), InstructBLIP Dai et al. (2024), MiniGPT-4 Zhu et al. (2023), Qwen-VL Bai et al. (2023), OpenFlamingo Awadalla et al. (2023), MMGPT Gong et al. (2023), and MiniGPT-v2 Chen et al. (2023) for zero-shot and few-shot (3-shot) inference. Additionally, LLaVA-1.5, InstructBLIP, and BLIP2 leverage supervised fine-tuning with QLoRA Dettmers et al. (2024).
  • CLIP-Based Methods: We include the original CLIP model as well as its extensions HateCLIPer and MOMENTA, which build upon CLIP’s contrastive embeddings to enhance hateful content detection.
For segmentation tasks, we utilize InstructBLIP, BLIP2, and CLIP+ITA (a version of HateSieve without pre-training). All baseline models are fine-tuned on the HatefulMemes dataset. Detailed segmentation procedures are provided in Appendix A.3.
Metrics
For HateSieve’s classification evaluation, we report Accuracy and F1-score, averaged over five independent runs. Evaluating the segmentation capabilities of HateSieve is challenging due to the absence of pixel-level and token-level annotations. To address this, we sampled 100 memes from the HatefulMemes dataset and conducted evaluations using both human annotators and LMMs Zheng et al. (2024) based on the following criteria:
  • Correctness: Determines whether the segmentation accurately captures the target social group or elements that reflect the hateful content, based on common-sense understanding.
  • Relevance: Assesses whether the highlighted image segments are meaningfully related to the highlighted text components, ensuring coherence between visual and textual elements.
Each criterion was scored using a binary system: 0 (No) or 1 (Yes). Implementation details for HateSieve and the LMM baselines are provided in Appendices A.4 and A.5, respectively. Comprehensive information on the segmentation evaluation process can be found in Appendix A.6.
4.2 Classification Results
Model HatefulMeme Harm-C Harm-P # t.p. $\downarrow$
Acc.$\uparrow$ F1$\uparrow$ Acc.$\uparrow$ F1$\uparrow$ Acc.$\uparrow$ F1$\uparrow$
@@ 4.2 Classification Results
[…]
Few-shot Learning
[…]
Supervised Fine-Tuning
InstructBLIP (13B) 63.55 59.34 65.54 42.52 51.98 36.68 65.72M
LLaVA-1.5 (13B) 66.34 63.28 61.61 56.88 59.57 58.62 65.72M
BLIP2 (6.7B) 62.85 56.43 54.28 55.68 45.91 41.37 33.35M
[…]
MOMENTA (434M) 73.34 70.02 83.82 82.80 89.84 88.26 7.73M
HateSieve (155M) 73.45 71.64 83.62 83.07 88.78 88.53 3.61M
Table 1: Model Performance Comparison. Bold scores indicate the best performance, while underlined scores denote the second-best performance. “Acc.” and “F1” represent classification accuracy and macro-F1 score, respectively. “# t.p.” denotes the number of trainable parameters.
Table 1 compares the classification performance of various LMMs and CLIP-based methods under zero-shot, few-shot, and supervised fine-tuning (SFT) settings. In the zero-shot scenario, GPT-4V clearly stands out among LMMs, achieving the highest accuracy (71.70%) and F1 score (71.28%) on the HatefulMemes dataset. By contrast, other open-source LMMs (e.g., CogVLM, LLaVA-1.5, and InstructBLIP) show limited capability, with lower accuracies (37.50%–65.20%) and F1 scores (27.28%–71.28%), revealing that pre-training alone is insufficient for capturing the nuanced semantics needed to detect hateful memes.
[…]
4.3 Segmentation Results
(a) Comparison of Segmentation Performance: $C$ represents the correctness score, while $R$ indicates the relevance score.
(b) Segmentation effect visualization.
Figure 5: Hateful content segmentation analysis.
Figure 5(a) demonstrates that HateSieve significantly outperforms InstructBLIP and BLIP2 in both correctness and relevance scores for segmentation, as evaluated by human annotators and LLM evaluators. In contrast, CLIP+ITA—which has not undergone pre-training—underperforms relative to the other models, underscoring the crucial role of contrastive learning pre-training in enhancing hateful content segmentation. Moreover, all models achieve slightly lower relevance scores compared to their correctness scores, suggesting that improvements are still needed to more accurately associate specific components within a hateful context. The inter-annotator agreement among human evaluators is discussed in Section A.9.
Figure 5(b) illustrates the segmentation results, supporting our observations from Figure 5(a). Specifically, CLIP+ITA without contrastive learning pre-training generates overly dispersed attention maps. While LMMs effectively identify relevant textual keywords through semantic reasoning, their image segmentation performance suggests that their classification capabilities for hateful memes rely more on the associated Large Language Models rather than on visual information.
@@ 5 Ablation Study
(a) Impact of ITA Parameter Scale and Triplet Data Scale on Model Performance.
CMGen Strategy HatefulMeme Harm-C Harm-P
Acc. F1 Acc. F1 Acc. F1
HateSieve 73.45 71.64 83.62 83.07 88.78 88.53
-w/o Inpainting 72.61 70.15 82.51 80.13 85.23 84.29
-w/ Text 71.28 69.43 81.79 81.05 86.38 84.06
(b) Impact of CMGen using different strategies.
(c) Impact of using different pre-training strategies.
Figure 6: Ablation studies on model performance.
@@
We examined how the number of self-attention layers within the Image-Text Alignment (ITA) module affects the classification performance of HateSieve. As shown in Figure 6(a), increasing the number of layers initially enhances classification accuracy. However, performance gains plateau and eventually decline when the layer count exceeds six, as evidenced by a noticeable decrease in the F1 score.
@@
We investigated the impact of varying the size of the triplet dataset used during the contrastive learning pre-training stage on the classification performance of HateSieve. As illustrated in Figure 6(a), we evaluated HateSieve’s performance when pre-trained with 0% (no pre-training), 25%, 50%, 75%, and 100% of the triplet dataset. The results demonstrate that increasing the amount of pre-training data consistently improves HateSieve’s classification capabilities.
CMGen Generation Strategy
We assessed how text captions influence the quality of the triplet dataset in the CMGen data generation process by evaluating: (1) the role of text captions in matching context-correlated memes based solely on images (-w/o inpainting), and (2) the impact of incorporating text embeddings when matching non-hateful pairs using FAISS (-w/ Text). As shown in Table 6(b), residual text captions impair classification performance, indicating interference with image information integration. Additionally, adding text embeddings to FAISS degraded triplet dataset quality, likely due to weak semantic correlations between meme text and images.
@@
We investigated how different pre-training strategies affect the classification performance of HateSieve by comparing three approaches: 1. In-domain Pre-training, utilizing only the HatefulMemes training set to directly sample negative image-text pairs without incorporating external reference datasets; 2. CLS-Only, replacing the contrastive learning pre-training task with a classification task using the triplet dataset; and 3. None, no pre-training. Our results in Figure 6(c) show that modifying the components of the triplet dataset or altering/removing the pre-training strategy negatively impacts model performance. Notably, adopting the CLS strategy resulted in a decline in performance on the Harm-C and Harm-P datasets that was as significant as no pre-training. This highlights that using classification as a pre-training task doesn’t ensure generalizability across various domains.
Transferability of Fine-Tuning Across Datasets
We evaluated the transferability of models fine-tuned on one dataset by testing them on different datasets. Table 2 summarizes the results, reporting both accuracy and F1 scores for three datasets: HatefulMemes (HM), Harm-C, and Harm-P.
Training Set HatefulMemes Harm-C Harm-P
Acc. F1 Acc. F1 Acc. F1
HM 73.45 71.64 72.54 69.32 72.13 70.02
Harm-C 65.29 63.82 83.62 83.07 80.54 78.52
Harm-P 63.73 61.28 76.44 73.26 88.78 88.53
Combined (All) 73.58 72.15 84.27 83.48 88.52 87.48
Table 2: Transferability of models fine-tuned on one dataset and tested on others. The best performance for each dataset is highlighted in bold.
Our findings indicate that models fine-tuned on HM perform best on the HM dataset, but their accuracy and F1 scores drop when evaluated on Harm-C and Harm-P. Conversely, models fine-tuned on Harm-C and Harm-P yield superior performance on their respective datasets, yet underperform on HM. This discrepancy can be attributed to concept drift: HM encompasses a broader range of hate speech categories (including racist, sexist, homophobic, and religious hate), whereas Harm-C and Harm-P predominantly feature memes related to COVID-19 and US politics.
Interestingly, fine-tuning on Harm-C and Harm-P results in less performance degradation when testing on each other’s datasets, suggesting that similar content domains exhibit lower concept drift. Moreover, combining all datasets for fine-tuning generally improves performance, highlighting the benefit of diverse and culturally varied training data to enhance generalizability.
Finally, even when fine-tuned on different datasets, our model consistently outperforms most LMMs in both zero-shot and QLoRA settings, demonstrating the robustness and effectiveness of our approach across various evaluation scenarios.
6 Category-Specific Evaluation
We conducted a manual inspection of 300 randomly selected memes from the HatefulMemes dataset, assigning each meme to a hate speech category based on consensus among three annotators. Table 3 presents the classification accuracy (Cls. Acc.), segmentation accuracy (Seg. Acc.), and consistency rate (Consis. Rate) for each category.
Category # Samples Cls. Acc. Seg. Acc. Consis. Rate
Racist 147 69.39 77.55 75.51
Sexist 45 86.67 100.00 86.67
Homophobic 12 50.00 100.00 50.00
Religion 78 73.08 76.92 65.38
Disability 21 71.43 100.00 71.43
Table 3: Results of a manual inspection of 300 memes. Cls. Acc. is the rate of correctly identifying memes as hateful or not, Seg. Acc. is the rate of correctly segmenting the target entities, and Consis. Rate is the proportion of cases where both classification and segmentation are either correct or incorrect.
While certain categories (e.g., Sexist, Disability) exhibit high segmentation accuracy, classification accuracy occasionally lags behind (e.g., Homophobic). This discrepancy suggests that locating offensive content does not always translate into correct classification. Categories like Racist and Religion display moderate performance in both metrics, highlighting the need for more diverse training data and targeted refinement. Overall, boosting classification consistency—particularly in underrepresented categories—remains an important goal for future work.
@@ 7 Conclusion
We developed HateSieve, a framework for classifying and segmenting hateful memes. Our experiments demonstrate that using contrastive learning with a custom triplet dataset enhances classification accuracy and achieves effective segmentation.
@@
Our work has several limitations that we plan to address in future research. First, our CMGen system primarily generates context-correlated memes based on image content rather than text, due to inherent restrictions (see Appendix A.7 for a detailed analysis). Second, achieving high accuracy in image segmentation within HateSieve remains challenging. Although our current approach uses attention maps at the image-patch level—and we have experimented with refining these maps to pixel-level detail via linear interpolation—this method introduces biases without substantially improving segmentation accuracy. Third, the current version of HateSieve focuses exclusively on English hate speech; we plan to extend support to additional languages in future releases. Finally, our framework is not specifically tailored to distinct social or cultural groups, largely due to the limited granularity of annotations in the existing dataset. Future work will concentrate on expanding dataset annotations and enhancing the system’s performance across a wider range of multimodal hate speech content.
@@ Appendix A Appendix
A.1 Segmentation Details
Figure 7: Hate component segmentation process with attention map.
A.2 Dataset
We utilize several datasets to train and to evaluate the performance of our HateSieve framework:
  • HatefulMemes Dataset Kiela et al. (2020): Provided by Facebook Research, this dataset comprises 10,000 annotated meme images that combine text and imagery. It is specifically designed to challenge models in detecting hate speech within memes by including subtle and multimodal instances of hateful content.
  • MMHS150k Dataset Gomez et al. (2020): This dataset contains 150,000 tweets, each paired with an image, collected between September 2018 and February 2019. The tweets were gathered using 51 Hatebase terms to explore hate speech on social media, offering a rich source of multimodal content for our study.
  • Harm-C and Harm-P Datasets Pramanick et al. (2021): Harm-C includes 3,544 memes focusing on COVID-19-related topics, while Harm-P comprises 3,552 memes related to U.S. politics. These datasets provide context-specific challenges for hate speech detection in memes.
By leveraging these diverse datasets, we aim to thoroughly evaluate our model’s ability to detect hateful content across different contexts and topics.
A.3 Segmentation with LMMs
To obtain both image and text segmentations using LMMs such as InstructBLIP and BLIP2, we employ the following prompt:Please examine the provided meme, which includes an [image] and accompanying [text]. Determine if the content can be considered hateful. If you conclude that the meme is hateful, identify and list the specific keywords or phrases in the text.
This prompt enables us to identify the text tokens that InstructBLIP considers ambiguous. For image segmentation, we adhere to the approach proposed by Li et al., which involves mapping the query corresponding to the Q-Former in InstructBLIP with the image’s cross-attention map using bilinear interpolation.
A.4 Implementation Details
Using the Contrastive Meme Generator, we produced a total of 42,344 triplet pairs. During the pre-training and fine-tuning phases, we employed the clip-vit-base-patch32 as our backbone for the image-text encoder and froze all the CLIP parameters. Our newly introduced Image-Text Alignment module comprises six layers of self-attention blocks. Additionally, we incorporated a two-layer MLP as a decoder for classification fine-tuning.
Figure 9: We compare the text and image embeddings of the original meme with its positive (Pos.) and negative (Neg.) pairs, as well as with other randomly selected images and texts (Oth.) from the same dataset, in the triplet dataset using Wasserstein, Euclidean, and Cosine distances. Lower distance values indicate higher similarity, providing a baseline for distance comparison.
A.5 LMMs Hyperparameters
@@
For supervised fine-tuning of LMMs, we adopted the QLoRA framework, incorporating trainable parameters ($d=64$) into the query and key components of the Q-Former. This modification was applied to joint LLM architectures, including OPT-6.7b for BLIP2 and Vicuna-7b for InstructBLIP, while keeping the original parameters frozen. We set a constant dropout rate of 0.05, fixed $\alpha$ at 256, and conducted fine-tuning with a learning rate of $5\times 10^{-5}$ and a batch size of 8.
A.6 Segmentation Evaluation
To assess our segmentation module, we employed both human and automated evaluations.
Human Evaluation:
Three independent annotators reviewed each segmented meme. Their evaluations were aggregated by majority vote to ensure reliability and minimize bias.
Automated Evaluation with LMMs:
We also used GPT-4V as an evaluator, tasking it to score the segmented memes using the same criteria as the human annotators. Both human evaluators and GPT-4V used the evaluation prompt illustrated in Figure 10.
A.7 Triplet Dataset Embedding Analysis
To verify that our CMGen produces context-correlated meme pairs, we conducted an analysis of text and image embedding distances with corresponding positive and negative pairs. We randomly selected 100 pairs from the triplet dataset. As shown in Figure 9, the image embedding distances for both positive and negative pairs (Wasserstein, Euclidean, and Cosine distances) are consistently lower than the baseline (Others) in most cases, indicating that CMGen successfully generates context-correlated images. However, the text embedding comparison shows that the distances are comparable to the baseline. This is largely because our current CMGen is primarily driven by images, and the text content often lacks detailed information, uses slang, or is challenging to mass-produce with LLMs due to safety policies. We aim to further enhance this aspect of CMGen in the future work.
A.8 Segmentation Results
Evaluation Prompt Given the following segmented meme image $\{I_{i}\}$ and accompanying text $\{T_{i}\}$ with highlighted tokens $[x_{i},\ldots,x_{j}]$, please evaluate the segmentation based on the criteria below. For each criterion, assign a score of 0 (No) or 1 (Yes) and provide a brief justification for your decision.
Correctness (Score: 0 or 1):
Does the segmentation accurately capture the target social group or elements that reflect the hateful content, based on common-sense understanding? Consider whether the highlighted areas in the image and text correspond to features commonly associated with the identified hateful content.
Relevance (Score: 0 or 1):
Are the highlighted image segments meaningfully related to the highlighted text components? Assess if the visual elements and the textual tokens work together to convey the intended message, especially in the context of the meme’s overall meaning. Please present your evaluation in the following format: Correctness: [Score]
Justification: [Your brief explanation]
Relevance: [Score]
Justification: [Your brief explanation]
Figure 10: Evaluation prompt provided to both human annotators and GPT-4V.
A.9 Inter-Annotator Agreement
We evaluated inter-annotator agreement among the three annotators by calculating Fleiss’ Kappa for both correctness and relevance. Table 4 presents the resulting values along with their interpretations.
Metric Fleiss’ Kappa Score Interpretation
Correctness 0.7572 Substantial Agreement
Relevance 0.6122 Moderate Agreement
Table 4: Fleiss’ Kappa scores for correctness and relevance.
These findings indicate that the annotators achieved substantial agreement on correctness and moderate agreement on relevance. Overall, the results underscore the reliability of our annotation process.